OpenStudy (anonymous):

Solve the equation 2 over 5 (x - 2) = 4x for x.

OpenStudy (unklerhaukus):

\[{2 \over 5 } (x - 2) = 4x \] multiply both sides of the equation by 5 , then divide both sides by 2

OpenStudy (unklerhaukus):

lastly add two to both side of the equation

OpenStudy (anonymous):

Do I multiply everything by 5?

OpenStudy (unklerhaukus):

\[{2 \over 5 } (x - 2) = 4x\] \[2+\frac12\left(5\times\left({2 \over 5 } (x - 2) \right)\right)=2+ \frac{5\times4x}2\]

OpenStudy (unklerhaukus):

yes

OpenStudy (jiteshmeghwal9):

\[\frac{2}{5(x-2)}=4x\]\[{2 \over 5x-10}={4x}\]cross multiply\[2=4x(5x-10)\]\[2=20x^2-40x\]\[0=20x^2-40x-2\]now solve it using quadratic formula i.e;\[-b \pm \sqrt{b^2-4ac}\over2a\]

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Blahhh Blahhhh((:

OpenStudy (anonymous):

what will you get using the quadratic formula?